how many molecules of tetraphosphorus decaoxide are present in 4.31 moles of this compound ?

and

how many moles of tetraphosphorus decaoxide are present in 1.07x10^22 molecules of this compound ?
Answer #1

We know that , 1 mole of substance contain Avogadro's number of particles.

\therefore 1 mole of P4O10 will contain 6.022 \times 10 23 P4O10 molecules.

\therefore P4O10 molecules present in 4.31 mol P4O10 = 4.31 mol P4O10\times (6.022 \times 10 23 P4O10 molecules /1 mol P4O10

P4O10 molecules present in 4.31 mol P4O10 = 2.595 \times 10 24 P4O10 molecules

ANSWER : 2.60 \times 10 24 P4O10 molecules

PART 2

mol of P4O10 present in 1.07 \times 10 22 P4O10 molecules

= 1.07 \times 10 22 P4O10 molecules \times ( 1 mol P4O10 / 6.022 \times 10 23 P4O10 molecules )

= 1.7768 \times 10 -03   mol P4O10

ANSWER : 1.78  \times 10 -03   mol P4O10